OpenStudy (anonymous):

wat is the solution to this equation? 12a=24

OpenStudy (anonymous):

12a/12 = 24/12 a = 2

OpenStudy (anonymous):

divid both sides by 12

OpenStudy (anonymous):

\[a=\frac{24}{12}=2\]

OpenStudy (anonymous):

12a=24 a=24/12=2